Sales $15,093 Depreciation expense 578 Tax expense 394 Interest expense, gross 171 Earnings from continuing operations (Net income) 685 EBITA 1,256 Cash 613 Average total assets 9,470 Total debt 2,908 Noncurrent deferred tax liabilities 391 Noncontrolling interest 0 Equity 871 Dividends paid 1,185 Cash from operating activities 2,451

how to calulate (round to one decimal place)

1. ebita to average assets

2.operating margin

3.debt to ebita

4.debt to book capitalization

5.retained cash flows to net debt
Answer #1

1. ebita to average assets = 1,256/9,470 = 13.3%

2.operating margin = 1,079/15,093 = 7.1%

3.debt to ebita = 2,908/1,834 = 1.6

4. debt to book capitalization = 2,908/4,170 = 69.7%

5.retained cash flows to net debt = 55.2%